Ask
your own question, for FREE!
Mathematics
14 Online
http://prntscr.com/6qim72
Still Need Help?
Join the QuestionCove community and study together with friends!
@TheSmartOne
are jar1 and jar2 independent?
yes
if they are, then the probability of drawing "A" from Jar A is 1/3, and the probability of drawing "white" from Jar B is 1/3, so the probability that one OR the other occurs, is the sum, or 1/3+1/3=2/3.
d'='
Still Need Help?
Join the QuestionCove community and study together with friends!
medal plz! ;)
ty ;)
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
4 hours ago
0 Replies
0 Medals
3 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
1 day ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal